16. Find the number of tangents to the curve y=-* passing through P = (1, 2).
Answer #1


Wi, (x,y) be point on the curve. y= 2 +) It must be on tagent line. let, (14 54,) be point on curve through which tangent pasAroup “ 念)。 》 31 = 0 → y = 2 * ) * - 2(1+x)* (1-x) (14 ) ? A: 24 42 +2.7-14 > T440 +1 = 0 3.%+ 2.25 +4 -3=0 (+2) - (a)=0 | (81. There are two (-2-v3, 4 + vz) through (1,2), tangent on the curre at & (-2+63 , $ - 42). passing